Four year ago the sum of the ages of mother and her two daughter was x years then after two years the sum of their age become........years
Answer by stanbon(75887)

Four year ago the sum of the ages of mother and her two daughter was x years then after two years the sum of their age become........years
----
Equation::
m-4 + a-4 + b-4 = x
Modify::
m+a+b = x - 12
----
Comment:: I assume that "after 2 yrs" means 2 years after 4 years ago.
That would be 2 years ago when the sum of their ages is
m-2+a-2+b-2 = m+a+b-6 = (x-12)+6 = x-6
--------------------------
Cheers,
Stan H.
------------------